# Quantum Pontus-Mpemba Effects in Real and Imaginary-time Dynamics

### v1

@@ Abstract
The quantum Pontus-Mpemba effect (QPME) is a counterintuitive phenomenon wherein a quantum system relaxes more rapidly through a two-step evolution protocol than through direct evolution under a symmetric Hamiltonian alone. In this protocol, the system first evolves under a symmetry-breaking Hamiltonian and then switches to a symmetric one. We demonstrate that QPME occurs under both real-time and imaginary-time dynamics with respect to $U(1)$-symmetry. Using tilted ferromagnetic initial states, we demonstrate that a transient asymmetric evolution significantly accelerates thermalization or convergence to the ground state for both real-time and imaginary-time evolutions, respectively. The effect is pronounced for small tilt angles, while larger tilts or antiferromagnetic initial states suppress it. Numerical evidence across different system sizes confirms the robustness of QPME, demonstrating its stability in the thermodynamic limit. This finding not only deepens our fundamental understanding of non-equilibrium quantum phenomena but also opens new possibilities for accelerating quantum simulation and computation.
@@ I Introduction
In recent years, research on the Mpemba effect has expanded beyond classical systems into the quantum realm, where it manifests as the quantum Mpemba effect (QME) [28, 29]. The QME can occur through two distinct ways: (1) in open quantum systems coupled to an external environment [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], or (2) in isolated quantum systems evolving under unitary dynamics. This work focuses on the latter case. This phenomenon has been observed in both integrable and chaotic quantum systems [42, 43, 44]. The QME manifests when the relaxation curves of interesting physical quantities cross one another for different initial conditions. For example, Refs [45, 46, 47, 48, 49, 50, 51, 52, 53, 54] show that $U(1)$-symmetry restoration occurs more rapidly for more asymmetric initial states under the $U(1)$-symmetric Hamiltonian quench. Likewise, such effect has also been explored in various contexts, including spin chains [55, 56, 57, 58, 59, 60, 61, 62, 63, 64], many-body localized systems [65], non-Hermitian systems [66, 67], and random quantum circuits [42, 68, 44]. QME also manifests in imaginary time evolution [69], where specific initial states of higher energy exhibit faster convergence to the ground state. This anomalous quantum phenomenon has gone beyond theoretical discussion, with experimental verification achieved in various quantum platforms [70, 71], marking a significant milestone in non-equilibrium quantum physics.
@@
In this work, we identify such non-equilibrium phenomena in both real-time and imaginary-time quantum evolution, with a specific focus on $U(1)$-symmetry. We term this phenomenon as quantum Pontus-Mpemba effect (QPME), distinguishing it from previous studies of the QME. Unlike the QME setup, where two systems $A$ and $B$ are prepared in different $U(1)$-asymmetric initial states and evolve under the same symmetric Hamiltonian $H_{sym}$, the QPME follows a distinct protocol. Here, systems $A$ and $B$ start in identical $U(1)$-asymmetric initial states. System $A$ evolves continuously under a $U(1)$-symmetric Hamiltonian $H_{sym}$. System $B$, however, first evolves under a $U(1)$-asymmetric Hamiltonian $H_{asym}$ for a controlled time interval before undergoing an abrupt transition to $H_{sym}$. Remarkably, the QPME manifests in two ways: (1) The time evolution of a given physical observable in system $B$ intersects that of system $A$ before both reach steady state, with $B$ exhibiting faster relaxation after the crossing, or (2) the evolution curve of system $B$ remains entirely below that of system $A$ at all times. Both cases demonstrate that the two-step protocol leads to accelerated relaxation compared to purely symmetric evolution. A comparison of the QME and QPME is presented in Table 1. Unlike the conventional QME, the QPME offers greater operational practicality for probing non-equilibrium quantum dynamics and accelerating quantum state preparation.
The remainder of this work is structured as follows. Section II introduces our theoretical framework, including the system setup and key observables. In Section III, we present comprehensive numerical results for both real-time and imaginary-time evolution, with a detailed analysis of the observed phenomena. Finally, Section IV provides concluding remarks and discusses potential implications of our findings.
@@ II Setup
$H=$ $-\sum_{j=1}^{L}(\sigma_{j}^{x}\sigma_{j+1}^{x}+\gamma\sigma_{j}^{y}\sigma_{j+1}^{y}+\mu\sigma_{j}^{z}\sigma_{j+1}^{z})$
@@
where $L$ is the total system size and $\sigma_{j}^{\alpha}$ ($\alpha=x,y,z$) are Pauli matrices acting on site $j$. The periodic boundary conditions enforce $\sigma_{L+1}^{\alpha}=\sigma_{1}^{\alpha}$ for $\alpha=x,y,z$. $\mu=-0.5$ sets the nearest-neighbor interaction strengths. Disorder is introduced via random fields $h_{j}$ along the $z$-axis, sampled uniformly from [W,W][-W,W] with $W=1$. In the Supplemental Material (SM), we provide evidence that the Hamiltonian $H$ is chaotic in that parameter regime. The parameter $\gamma$ serves as a control for $U(1)$-symmetry breaking in the system: at $\gamma=1$, the Hamiltonian preserves $U(1)$-symmetry and is denoted $H_{sym}$, while any deviation from unity explicitly breaks this symmetry, yielding the asymmetric Hamiltonian $H_{asym}$.
@@
Figure 3: (a) Imaginary-time evolution of $\langle\hat{H}\rangle_{\tau}$ for tilted ferromagnetic states ($\theta=0.05\pi$) at system sizes $L=12$ (light colors) and $L=48$ (dark colors), with red, blue, and green curves maintaining have the same meaning as Fig. LABEL:fig:_E_imag_time_F. (b) Energy density difference between two protocols as a function of $\tau$ for both $L=12$ and $L=48$. $\langle\hat{H}\rangle_{\tau}^{1}$ and $\langle\hat{H}\rangle_{\tau}^{2}$ represent the energy expectation values calculated from direct quench and two-step protocol evolution, respectively.
@@ III.1 Real-time Dynamics
We investigate the dynamics of entanglement asymmetry by considering a subsystem $A$ of 4 sites and setting the switching time from $H_{asym}$ to $H_{sym}$ at $t=0.5$. Our analysis compares a direct relaxation under $H_{sym}$ with our two-step protocol, revealing distinct behaviors for different initial states. For a tilted ferromagnetic initial state with a tilt angle of $\theta=0.2\pi$, we observe the clear emergence of case (1) of QPME, as shown in Fig. 1(a). This is identified by the crossing of the EA curves for the direct quench (red) and the two-step protocol (green) around $t=1$. The underlying mechanism is the phenomenon of Hilbert subspace imprint (HSI) [79], where certain initial states have significant overlap with only a small, non-ergodic subset of the Hilbert space, leading to non-thermalization dynamics. To understand why the direct quench is slow, we note that the dynamics of the tilted ferromagnetic state under $H_{sym}$ are unitarily equivalent to the dynamics of the ferromagnetic state under a different, corresponding weak asymmetric Hamiltonian $H^{\prime}$ (see SM for details). In this equivalent picture, the symmetry-breaking terms in $H^{\prime}$ are proportional to $\sin{\theta}$ and act as a weak perturbation for small tilt angles. According to HSI, such weak symmetry-breaking results in the ferromagnetic state having a limited overlap with the full spectrum of eigenstates, causing slow, non-thermal relaxation. The two-step protocol overcomes this limitation as the initial evolution under the symmetry-breaking Hamiltonian enhances the thermal properties of the state by promoting its spread across the Hilbert space. This process destroys the condition for HSI and accelerates the subsequent relaxation when the system is switched to $H_{sym}$, leading to the observed QPME. We further observe that the green and red curves converge to different steady-state values, despite evolving under the same final Hamiltonian after the switch. This difference arises because direct evolution under $H_{sym}$ alone results in a non-thermal steady state due to HSI, whereas the two-step protocol drives the system toward a thermal state. However, the QPME vanishes for larger tilt angles, such as $\theta=0.4\pi$ (see SM). At a large tilt, the initial state is already distributed broadly across different charge sectors with no remnants of HSI. This leaves no room for the transient asymmetric evolution to further enhance thermalization, thereby eliminating the speed-up effect. Furthermore, we note that the transient asymmetric evolution induces a characteristic overshooting in the EA (green curve), a feature that has been previously reported in [62] for symmetry-breaking dynamics.
@@
In stark contrast, QPME is absent for the tilted antiferromagnetic initial states, as demonstrated for angles $\theta=0.2\pi$ and $\theta=0.4\pi$ in Fig. 1(b) and SM. Here, the EA for the two-step protocol remains consistently above that of the direct quench. This is because the antiferromagnetic state already resides in the $Q=0$ sector, which has the largest Hilbert space dimension. Unlike the ferromagnetic state, it is inherently more “thermal” and does not suffer from HSI. Consequently, the initial asymmetric evolution provides no additional thermalization advantage, and no QPME is observed. As the initial state is already “thermal” in nature, both protocols ultimately converge to the same steady-state EA value.
@@ III.2 Imaginary-time Dynamics
To verify that the observed QPME is a robust physical phenomenon, we systematically investigate its dependence on system size. We compare energy dynamics for a system of size $L=12$, computed via exact methods, with a significantly large system of $L=48$, simulated using time-evolving block decimation (TEBD) [80]. As expected for an extensive quantity, the total energy in the $L=48$ system is four times larger than in the smaller system (Fig. 3(a)). To directly quantify the QPME, we analyze the energy density difference between the two protocols. The result, presented in Fig. 3(b), demonstrates that the difference curves for $L=12$ and $L=48$ are nearly identical. This confirms that the relative speed-up from the two-step protocol is independent of system size. This scaling behavior suggests that QPME is a genuine feature of the system’s dynamics and is expected to survive in the thermodynamic limit. We also investigate the effect of different switching times and find that the QPME is most pronounced when the switching time is significantly shorter than the system’s intrinsic relaxation timescale (see SM).
@@
IV Discussions and Conclusion

### v2

@@ Abstract
The quantum Pontus-Mpemba effect (QPME) is a counterintuitive phenomenon wherein a quantum system relaxes more rapidly through a two-step evolution protocol than through direct evolution under a symmetric Hamiltonian alone. In this protocol, the system first evolves under a symmetry-breaking Hamiltonian and then switches to a symmetric one. We demonstrate that QPME occurs under both real-time and imaginary-time dynamics with respect to $U(1)$-symmetry. Using tilted ferromagnetic initial states, we demonstrate that a transient asymmetric evolution significantly accelerates thermalization or convergence to the ground state for both real-time and imaginary-time evolutions, respectively. The effect is pronounced for small tilt angles, while larger tilts or antiferromagnetic initial states suppress it. Further numerical evidence confirms the consistency of QPME across different system sizes studied. Finally, by implementing a variational optimization procedure, we identify an optimized dynamical path, demonstrating that the relaxation rate can be further enhanced through tailored symmetry-breaking protocols. This work extends the framework of nonequilibrium quantum phenomena to incorporate active state preparation, with direct implications for the implementation of quantum simulation.
@@ I Introduction
In recent years, research on the Mpemba effect has expanded beyond classical systems into the quantum realm, where it manifests as the quantum Mpemba effect (QME) [5, 93]. The QME can occur through two distinct ways: (1) in open quantum systems coupled to an external environment [67, 24, 95, 63, 1, 52, 87, 19, 38, 71, 31, 57, 62, 26, 27], or (2) in isolated quantum systems evolving under unitary dynamics. This work focuses on the latter case. This phenomenon has been observed in both integrable and chaotic quantum systems [7, 59, 80]. The QME manifests when the relaxation curves of interesting physical quantities cross one another for different initial conditions. For example, Refs [35, 34, 32, 36, 16, 82, 20, 70, 14, 8, 76] show that $U(1)$-symmetry restoration occurs more rapidly for more asymmetric initial states under the $U(1)$-symmetric Hamiltonian quench. Likewise, such effect has also been explored in various contexts, including spin chains [49, 73, 65, 37, 9, 89, 94, 92, 17, 72], many-body localized systems [58], non-Hermitian systems [42, 30], Sachdev-Ye-Kitaev models [86, 21] and random quantum circuits [7, 51, 80]. QME also manifests in imaginary time evolution [25], where specific initial states of higher energy exhibit faster convergence to the ground state. This anomalous quantum phenomenon has gone beyond theoretical discussion, with experimental verification achieved in various quantum platforms [46, 88], marking a significant milestone in non-equilibrium quantum physics.
@@
In this work, we identify such non-equilibrium phenomena in both real-time and imaginary-time quantum evolution, with a specific focus on $U(1)$-symmetry. We term this phenomenon as quantum Pontus-Mpemba effect (QPME), distinguishing it from previous studies of the QME. Unlike the QME setup, where two systems $A$ and $B$ are prepared in different $U(1)$-asymmetric initial states and evolve under the same symmetric Hamiltonian $H_{sym}$, the QPME identifies an optimal dynamical pathway for a single initial state via a tailored two-step protocol. Here, systems $A$ and $B$ start in identical $U(1)$-asymmetric initial states. System $A$ evolves continuously under a $U(1)$-symmetric Hamiltonian $H_{sym}$. System $B$, however, first evolves under a $U(1)$-asymmetric Hamiltonian $H_{asym}$ for a controlled time interval before undergoing an abrupt transition to $H_{sym}$. Remarkably, the QPME manifests in two ways: (1) The time evolution of a given physical observable in system $B$ intersects that of system $A$ before both reach steady state, with $B$ exhibiting faster relaxation after the crossing, or (2) the evolution curve of system $B$ remains entirely below that of system $A$ at all times. Both cases demonstrate that the two-step protocol leads to accelerated relaxation compared to purely symmetric evolution. A comparison of the QME and QPME is presented in Table 1. This operational distinction makes the QPME a more practical and active strategy for accelerating quantum state preparation and probing non-equilibrium dynamics in settings where one seeks an efficient path from a given initial condition to a target state, rather than comparing different starting points.
The remainder of this work is structured as follows. Section II introduces our theoretical framework, including the system setup and key observables. In Section III, we present comprehensive numerical results for both real-time and imaginary-time evolution, with a detailed analysis of the observed phenomena. In Section IV, we employ a variational optimization approach to identify an optimized dynamical path towards equilibrium. Finally, Section V provides concluding remarks and discusses potential implications of our findings.
@@ II Setup
$H=$ $-\sum_{j=1}^{L}(\sigma_{j}^{x}\sigma_{j+1}^{x}+\gamma\sigma_{j}^{y}\sigma_{j+1}^{y}+\mu\sigma_{j}^{z}\sigma_{j+1}^{z})$ (3)
@@
where $L$ is the total system size and $\sigma_{j}^{\alpha}$ ($\alpha=x,y,z$) are Pauli matrices acting on site $j$. The periodic boundary conditions enforce $\sigma_{L+1}^{\alpha}=\sigma_{1}^{\alpha}$ for $\alpha=x,y,z$. $\mu=-0.5$ sets the nearest-neighbor interaction strengths. Disorder is introduced via random fields $h_{j}$ along the $z$-axis, sampled uniformly from [W,W][-W,W] with $W=1$. In the Supplemental Material (SM), we calculate the level spacing ratio [68] within the half-filling sector to ensure that the Hamiltonian $H$ is chaotic in that parameter regime. The parameter $\gamma$ serves as a control for $U(1)$-symmetry breaking in the system: at $\gamma=1$, the Hamiltonian preserves $U(1)$-symmetry and is denoted $H_{sym}$, while any deviation from unity explicitly breaks this symmetry, yielding the asymmetric Hamiltonian $H_{asym}$.
@@
Figure 3: (a) Imaginary-time evolution of $\langle\hat{H}\rangle_{\tau}$ for tilted ferromagnetic states ($\theta=0.05\pi$) at system sizes $L=12$ (light colors) and $L=48$ (dark colors), with red, blue, and green curves maintaining have the same meaning as Fig. 2. (b) Energy density difference between two protocols as a function of $\tau$ for both $L=12$ and $L=48$. $\langle\hat{H}\rangle_{\tau}^{1}$ and $\langle\hat{H}\rangle_{\tau}^{2}$ represent the energy expectation values calculated from direct quench and two-step protocol evolution, respectively.
@@ III.1 Real-time Dynamics
We investigate the dynamics of entanglement asymmetry by considering a subsystem $A$ of 4 sites and setting the switching time from $H_{asym}$ to $H_{sym}$ at $t=0.5$. Our analysis compares a direct relaxation under $H_{sym}$ with our two-step protocol, revealing distinct behaviors for different initial states. For a tilted ferromagnetic initial state with a tilt angle of $\theta=0.2\pi$, we observe the clear emergence of case (1) of QPME, as shown in Fig. 1(a). This is identified by the crossing of the EA curves for the direct quench (red) and the two-step protocol (green) around $t=1$. The underlying mechanism is the phenomenon of Hilbert subspace imprint (HSI) [91]. This phenomenon occurs when a system, initially prepared in a specific symmetry sector of a symmetric Hamiltonian $H_{0}$, is subjected to a weak symmetry breaking perturbation $H_{1}$. In this scenario, the initial state’s overlap remains predominantly confined to a polynomially small set of eigenstates of the perturbed Hamiltonian $H=H_{0}+H_{1}$, as these eigenstates maintain an adiabatic connection to those of $H_{0}$. This dynamical confinement thereby leads to non-thermalizing behavior. To understand why the direct quench is slow, we note that the dynamics of the tilted ferromagnetic state under $H_{sym}$ are unitarily equivalent to the dynamics of the ferromagnetic state under a different, corresponding weak asymmetric Hamiltonian $H^{\prime}$ (see SM for details). In this equivalent picture, the symmetry-breaking terms in $H^{\prime}$ are proportional to $\sin{\theta}$ and act as a weak perturbation for small tilt angles. According to HSI, such weak symmetry breaking approximately confines the dynamics of the ferromagnetic state to a polynomially small corner of the Hilbert space, causing slow, non-thermal relaxation. The two-step protocol overcomes this limitation as the initial evolution under the symmetry-breaking Hamiltonian enhances the thermal properties of the state by driving the state into a broader charge distribution and promoting its spread across the Hilbert space. This process destroys the condition for HSI and accelerates the subsequent relaxation when the system is switched to $H_{sym}$, leading to the observed QPME. We further observe that the green and red curves converge to different steady-state values, despite evolving under the same final Hamiltonian after the switch. This difference arises because direct evolution under $H_{sym}$ alone results in a non-thermal steady state due to HSI, whereas the two-step protocol drives the system toward a thermal state. However, the QPME vanishes for larger tilt angles, such as $\theta=0.4\pi$ (see SM). At a large tilt, the initial state is already distributed broadly across different charge sectors with no remnants of HSI. This leaves no room for the transient asymmetric evolution to further enhance thermalization, thereby eliminating the speed-up effect. Furthermore, we note that the transient asymmetric evolution induces a characteristic overshooting in the EA (green curve), a feature that has been previously reported in [92] for symmetry-breaking dynamics.
@@
In stark contrast, QPME is absent for the tilted antiferromagnetic initial states, as demonstrated for angles $\theta=0.2\pi$ and $\theta=0.4\pi$ in Fig. 1(b) and SM. Here, the EA for the two-step protocol remains consistently above that of the direct quench.This absence stems from the intrinsic properties of the antiferromagnetic state: it initially possesses a high overlap with an exponential number of eigenstates within the vast $Q=0$ sector. It is therefore inherently “thermal”, meaning it efficiently explores the available state space and relaxes rapidly under $H_{\text{sym}}$ alone, without suffering from the HSI confinement that plagues the ferromagnetic case. Consequently, the initial asymmetric evolution provides no additional thermalization advantage, and no QPME is observed. As the initial state is already “thermal” in nature, both protocols ultimately converge to the same steady-state EA value.
@@ III.2 Imaginary-time Dynamics
To investigate the persistence of the QPME, we compare energy dynamics for a system of size $L=12$, computed via exact methods, with a significantly large system of $L=48$, simulated using time-evolving block decimation (TEBD) [81]. As expected for an extensive quantity, the total energy in the $L=48$ system is four times larger than in the smaller system (Fig. 3(a)). To directly quantify the QPME, we analyze the energy density difference between the two protocols. The result, presented in Fig. 3(b), demonstrates that the difference curves for $L=12$ and $L=48$ are nearly identical. This observed consistency across system sizes indicates that the relative speed-up from the two-step protocol is not a finite-size artifact and suggests its manifestation in larger systems. We also investigate the effect of different switching times and find that the QPME is most pronounced when the switching time is significantly shorter than the system’s intrinsic relaxation timescale (see SM).
@@
IV Variational optimization of the dynamical path
Figure 6: Variational identification of the relaxation path for $L=12$ and tilted ferromagentic state with $\theta=0.2\pi$. (a) Convergence of the loss function (energy at $\tau_{total}=0.4$) over optimization steps. The switching time is fixed at $\tau=0.1$. The black dashed line indicates the reference energy for a system governed solely by $H_{sym}$. (b) Energy relaxation during imaginary-time evolution. The red line shows the standard direct quench under $H_{sym}$. The two-step protocols involve an initial evolution under $H_{opt}$ followed by $H_{sym}$. The blue line utilizes the optimized parameters $\{\gamma_{j},h_{j}\}$, while the green dashed line uses unoptimized parameters. The inset displays the energy difference between the direct quench $\langle\hat{H}\rangle_{\tau}^{1}$, and the two-step protocol $\langle\hat{H}\rangle_{\tau}^{2}$ as a function of $\tau$. A positive (negative) difference indicates that the energy expectation value of the two-step protocol relaxes faster (slower) than that of the direct quench.
In the previous sections, we demonstrated the QPME using a uniform symmetry-breaking Hamiltonian $H_{asym}$ with a fixed parameter $\gamma$. A natural question arises: can we identify a theoretically optimized path that enhances the relaxation rate? To rigorously identify a highly efficient dynamical path towards equilibrium without relying on an arbitrary ansatz, we implement a variational optimization procedure over the parameters of the transient Hamiltonian. We parameterize the transient Hamiltonian $H_{opt}$ using the same form as Eq. (3) but relax the constraints to allow for site-dependent couplings and fields. This yields a high-dimensional variational landscape defined by:
$H_{opt}(\{\gamma_{j}\},\{h_{j}\})=$ $-\sum_{j=1}^{L}(\sigma_{j}^{x}\sigma_{j+1}^{x}+\mu\sigma_{j}^{z}\sigma_{j+1}^{z})$ (9)
$-\sum_{j=1}^{L}\gamma_{j}\sigma_{j}^{y}\sigma_{j+1}^{y}+\sum_{j=1}^{L}h_{j}\sigma_{j}^{z}.$
Here, the variational parameters are the bond-dependent anisotropy $\gamma_{j}\in[0,1]$, which controls the strength of the symmetry-breaking term on each bond, and the site-dependent longitudinal fields $h_{j}\in[-1,1]$. This parameterization covers a wide class of local symmetry-breaking Hamiltonians.
We define the efficiency of the dynamical path by the proximity of the system to the target ground state after a fixed total evolution time, $\tau_{total}$. The optimization objective is to minimize the expectation value of the energy with respect to the target symmetric Hamiltonian $H_{sym}$ (where $\gamma_{j}=1$ for all $j$):
$\mathcal{L}=\langle\psi(\tau_{total})|\hat{H}_{sym}|\psi(\tau_{total})\rangle$ (10)
where $|\psi(\tau_{total})\rangle$ is the state obtained by evolving the initial state first under $H_{opt}(\{\gamma_{j}\},\{h_{j}\})$ for time $\tau$, and subsequently under $H_{sym}$ for the remainder of the time. In the optimization procedure, we enforce physical constraints on the variational parameters to ensure that $\gamma_{j}\in[0,1]$ and $h_{j}\in[-1,1]$ throughout the search. This is implemented via a projected gradient-descent method, where after each update step, any parameter exceeding its prescribed bounds is clipped back to the nearest boundary. Such regularization guarantees that the transient Hamiltonian remains physically admissible and allows for a meaningful comparison with the unoptimized, uniform perturbation case. To compare with the unoptimized results from Fig. 2(c), we perform this optimization for the imaginary-time dynamics of a tilted ferromagnetic state ($\theta=0.2\pi$) with a system size $L=12$. We fix the total evolution time $\tau_{total}=0.4$ and the switching time $\tau=0.1$. The optimizer is initialized with uniform couplings $\gamma_{j}=0.2$ for all $j$ and random longitudinal fields, using a learning rate of $0.4$.
As shown in Fig. 6(a), the loss function (final energy) decreases monotonically as the optimization progresses and saturates after approximately 8 steps. This indicates that the gradient descent algorithm consistently converges to a specific, non-trivial configuration of couplings $\{\gamma_{j}\}$ and fields $\{h_{j}\}$. To verify that this configuration indeed yields a faster relaxation rate, we analyze the time dependence of the energy in Fig. 6(b). For the two-step quench using unoptimized parameters (green dashed line), we observe the first scenario of QMPE described in Table. 1. This is characterized by a crossing between the two-step evolution and the direct quench (red line) around $\tau=0.32$, implying the advantage only emerges at later times. In contrast, when employing the optimized parameters for $H_{opt}$ (blue line), we observe another case of QMPE behavior, where the energy remains strictly below the direct quench baseline throughout the entire evolution. This persistent advantage is quantified in the inset, where the energy difference between the direct quench and the optimized protocol remains positive for all $\tau$.
We also extend this variational approach to real-time dynamics to minimize the entanglement asymmetry, with detailed results provided in the SM. In doing so, it is important to address a subtle distinction regarding the definition of an optimized paths. In imaginary-time evolution, the system universally converges to the unique ground state of $H_{sym}$, ensuring an identical target for both direct and two-step protocols. However, in real-time dynamics, the system thermalizes to a steady state determined by its energy density. Since the initial evolution under the variational Hamiltonian $H_{opt}$ modifies the system’s energy relative to $H_{sym}$, the two-step protocol ultimately targets a thermal ensemble with a different effective temperature than the direct quench. Nevertheless, the entanglement asymmetry eventually reaches to zero for both protocols. Consequently, the optimization remains physically meaningful, as it aims to maximize the rate of symmetry restoration. This highlights a key conceptual difference between the two regimes: while imaginary-time optimization targets a unique ground state, real-time optimization focuses on accelerating the dynamical process of symmetry recovery.
These results demonstrate that the optimized configuration generates a trajectory with a significantly faster decay rate than the unoptimized perturbation. This confirms that our variational method successfully identifies a uniquely efficient dynamical path toward equilibrium within the variational landscape. Therefore, the principle of the Pontus-Mpemba effect, integrated with variational optimization, provides practical engineering guidance for fast state preparation.
V Discussions and Conclusion